Circular No. 9059

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 O4 (HILL) R. E. Hill reports his discovery of another comet on CCD images taken

with the Catalina 0.68-m Schmidt telescope (discovery observation tabulated below); he notes that four stacked 20-s exposures in 2′′ seeing show the object to have a bright, nuclear condensation of size 15′′ × 20′′ that is elongated northwest-southeast, with no coma and a broad, diffuse tail ∼ 2′ long in p.a. 300^{o}. After posting on the Minor Planet Center’s ‘NEOCP’ webpage, other cometary astrometrists have noted the object’s cometary appearance. Y. Ikari (Moriyama, Shiga-Ken, Japan, 0.26-m reflector, July 30.7 UT) reports coma diameter 0′.3 and a tail 0′.6 long in p.a. 315^{o}. F. Fratev (Zvezdno Obshtestvo Observatory, Plana, Bulgaria, 0.35-m reflector, July 31.0) writes that sixty stacked 40-s unfiltered exposures reveal a coma nearly 28′′ in diameter and a tail ∼ 1′.5 long in p.a. 320^{o}. J. A. Henriquez (Tenerife, Spain, 0.2-m reflector, July 31.1; poor weather conditions) measures an 11′′ coma and a 14′′.5 tail in p.a. 317^{o}. J. E. McGaha (Tucson, AZ, U.S.A., 0.36-m reflector, July 31.3) notes that six stacked 60-s images show an 8′′ bright round coma with a fan-shaped tail 35′′ long at p.a. 315^{o}. W. H. Ryan (Magdalena Ridge 2.4-m reflector, July 31.35–31.38) notes a tail at p.a. ∼ 290^{o}. G. Hug (Scranton, KS, U.S.A., 0.56-m reflector, July 31.4) reports a short, broad tail in p.a. ∼ 310^{o}.

2009 UT α_{2000} δ_{2000} Mag. Observer
July 30.37424 0^{h}33^{m}09.^{s}91 +7^{o}38′01′′.8 16.3 Hill

The available astrometry, the following preliminary parabolic orbital elements by B. G. Marsden, and an ephemeris appear on MPEC 2009-O56.

T = 2009 Dec. 1.451 TT  ω = 210.^{o}640
Ω = 173.459 }2000.0
q = 2.86118 AU  i = 94.583

COMETS C/2009 D5 AND C/2009 D6 (SOHO) Further to IAUC 9058, starlike apparent comets have been found on

SOHO website images: C/2009 D5 was a Kreutz object (mag ∼ 6.5–7), while Meyer-group member C/2009 D6 was slightly elongated (mag ∼ 7).

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 D5 Feb. 23.679 22^{h}48.^{m}1 −10^{o}41′ C3 MK 2009-M55
C/2009 D6  24.814 22 37.2 − 7 26 C2 RK 2009-M55
2009 July 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green